Circular No. 8242

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003kc IN MCG +05-23-37 Further to IAUC 8241, J. Burket and W. Li report the LOSS discovery,

on unfiltered KAIT images taken on Nov. 21.6 (mag ∼ 18.8) and 22.6 UT (mag ∼ 18.1), of a supernova located at α = 9^{h}46^{m}34^{s}.32, δ = +30^{o}39′19′′.0 (equinox 2000.0), which is 5′′.4 west and 8′′.2 north of the nucleus of MCG +05-23-37. A KAIT image taken on Apr. 2.2 showed nothing at this position (limiting mag ∼ 19.5).

R. J. Foley, B. J. Swift, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Nov. 22 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003kc is of type Ia, similar to SN 1991T (Filippenko et al. 1992, Ap.J. 384, L15). Strong Fe III features suggest that it is near maximum brightness.

SUPERNOVAE 2003jz, 2003ka, AND 2003kb Foley et al. add that inspection of CCD spectra, obtained as above on

Nov. 22 UT, reveals that SN 2003kb (IAUC 8241) is probably of type Ia, ∼ 2 weeks past maximum brightness.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, write that a spectrum (range 370–750 nm) of SN 2003jz (cf. IAUC 8241), obtained by L. Macri on Nov. 20.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova a few weeks past maximum. A spectrum of SN 2003ka (cf. IAUC 8241), also obtained by Macri on Nov. 20.50, shows it to be a peculiar type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 5761 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 8300 km/s.

COMET C/2003 T4 (LINEAR) Improved parabolic orbital elements from MPEC 2003-W12:

T = 2005 Apr. 3.1380 TT  ω = 181.^{o}8410
Ω = 93.8487 }2000.0
q = 0.844473 AU  i = 86.7148
2003 November 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green